# Through the Lens of Human-Human Collaboration: A Configurable Research Platform for Exploring Human-Agent Collaboration

### v1

@@ 3.2.1. Participant Interface
Figure 3. An example of the participant interface for the Shape Factory experiment (Bos et al., 2004). The interface consists of five fully configurable modules: My Status displays participant’s private status, My Actions displays permitted actions, My Tasks displays the tasks to be completed, Social provides available interaction channels, and Information Dashboard shows other participants’ real-time information. The timer of the experiment is shown at the top.
The participant interface uses a three-column layout that generalizes across common experimental paradigms by separating visual modules for private state, social interaction, task operations, and shared awareness. An example participant interface for the Shape Factory experiment, which is used in the case study in Section 4.1, is shown in Figure 3.
The left column presents private state and individual actions. The My Status module shows individual resources and role-dependent capacities that are not necessarily visible to others, which could support studies of interdependence and strategic planning. The My Action module provides affordances for private actions, such as producing a resource or making an independent decision. The My Task module displays individual or collaborative goals and fulfillment status, which could support shared plans and reduce coordination overhead by making the task structure explicit.
The middle column contains the Social module for all participant interactions. This is supported by fruitful CSCW research that has underscored the importance of communication and negotiation affordances for studying trust, information grounding, and collaboration effectiveness (Straus, 1997; Bos et al., 2002; Hinds and Bailey, 2003). Here, a Trading page supports resource exchange and can be adaptive to broader collaborative decision-making activities. A separate Chat page supports text-based communication between participants. Researchers can configure trading with free-form or protocol-bounded offers to study negotiation dynamics and coalition formation (Poole et al., 1991). For the chat functionalities, we implemented three variations: private messages, group messages, or be disabled entirely. LLM agents can interact with human participants through trading and messaging by following the instructions and action spaces specified in ACP. The right column is a fully configurable shared Information Dashboard. The dashboard makes workspace awareness (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002) a customizable visual module; researchers can select what information to disclose, to whom, and at what level of granularity.
@@ 3.2.2. Researcher Interface
The researcher interface provides direct controls for researchers to design, execute, monitor, and analyze experiments for human-agent collaboration research. The final design, as shown in Fig 4, was refined through a participatory cognitive walkthrough (Polson et al., 1992) with five HCI researchers in human-AI collaboration to solicit design and interaction feedback and align the workflow with their needs (details in Section 4.2). The researcher interface organizes functionalities into three pages: experiment setup, real-time monitoring, and post-study analysis. For experiment setup, the interface guides researchers through four primary steps: experiment selection, parameter configuration, interaction control, and participant registration, as shown in Figure 4. In addition, researchers can customize participants’ information, including the persona profile for LLM agents, without editing the code.
The Experiment Selection module allows a researcher to load a saved session or create a new one. Once the session is created, it provides a list of experimental paradigms, each with a text description and a simplified illustration, so that researchers can become familiar with the purpose of different experimental paradigms, even if they do not know any of the experiments beforehand. Researchers can also upload a custom paradigm defined in an ECL file.
Then, the Session Configuration module offers researchers comprehensive control over (1) experiment parameters, such as initial resource allocation and experiment constraints; (2) interaction controls, such as communication channels and information dashboard (details in Section 3.3); (3) agent design, including parameters like response latency; (4) participant registration for both humans and agents.
Once the session is live, the Real-time Monitoring module allows the researcher to observe participant interactions and progress. After the session is completed, the Result Analysis module provides summary statistics and visualizations of participant performance and behavior. Researchers can view aggregate results or filter by participant, and can export raw event logs for further analysis.
Figure 4. The workflow of HCI researchers for study setup through the researcher interface, which includes five steps: experiment selection, parameter configuration, interaction control, participant registration, and participant customization (optional).
@@
An overview of HCI researchers’ workflow is illustrated in Figure 4. For the study setup, researchers begin by selecting whether they want to load an existing session or a new session; if the latter, they first need to choose the experimental paradigms that best suit their research interests. The system also enables researchers to upload a customized experiment configuration template, using the language specified in Section 3.2.3, if the provided ones do not meet their needs. Then, researchers will proceed to the session configuration module to customize interaction controls and agent-specific designs. The horizontal scrolling function of the researcher interface allows researchers to iteratively adjust the parameters in accordance with the experimental procedures or manipulate the interaction controls with different sets of parameters, which is highlighted by participants in the cognitive walkthrough study. After completing all the configurations, researchers conclude the setup with participant registration to decide the number of human and agent participants involved.
@@ 3.2.3. Experiment Configuration Language (ECL)
The ECL is specified in a structured format by defining its components through a set of core primitives: objects, actions, policies, and views. The language is organized around these four fundamental concepts. Objects represent the primary entities within the experiment, such as resources (e.g., money, parts), roles (e.g., buyer, seller), or abstract concepts like team scores. Each object is defined by a set of attributes, such as initial_valueinitial\_value or visibilityvisibility. Actions define the operations available to participants, specifying the actor, the costs an action incurs, and its effects on the state of the system’s objects. Policies are logical constraints that are specified by different experimental paradigms. These range from preconditions for specific actions (e.g., requiring sufficient resources) to global rules like time limits or winning conditions. Finally, Views separates the system’s internal state from the presentation layer by defining which objects and attributes are rendered in specific modules of the participant interface. This separation is critical, as it allows researchers to manipulate information visibility without altering the experiment’s core logic.
A concrete example of ECL for the Shape Factory is provided in Appendix B. In this example, $Money$ and $Shape$ are defined as objectsobjects. the $ProduceAction$ links the costs and effects for producing a shape, and is governed by a policy that validates the actor’s resources. The $Views$ declares what objects and states are visible in each visual module. During setup, the Experiment Controller parses an ECL file to dynamically instantiate the environment, enforce the specified policies, and render the correct views for each participant. The experiment controller also validates the configurations and reports conflicts when a new configuration template is uploaded. This approach allows researchers to create, share, and modify experimental paradigm designs in a transparent and reproducible manner.
@@ 3.2.4. Agent Context Protocol
Inspired by the emerging standard of model context protocol [^3] https://modelcontextprotocol.io/ that connects LLMs with external tools and resources, we designed the ACP to standardize communication between LLM agents and the experiment environment (Figure 5). ACP aims to establish a standardized channel for perception, information transfer, and interaction between LLM agents and the experiment environment to ensure robust agent integration regardless of agent architecture designs. As a result, ACP enforces that LLM agents’ behaviors are aligned with the experimental rules and valid action spaces. More importantly, agents perceive the experiment and act within it under the same constraints and through the same structured channels as a remote human participant. This parity is essential for the validity of controlled, comparative studies.
At initialization, the experiment controller (Section 3.2.5) constructs an experiment-specific context for agents from the experiment configurations and the interaction conditions specified by researchers. The context includes participant IDs, groups, private and public states, permitted actions with schemas, communication channels, perception interval, and experiment rules. The perception interval is a time parameter for determining how often the experiment controller issues periodic state summaries of the current situation to the agents. The experiment controller will validate that the information in the state summaries is aligned with the private and public states specified in the agent context.
During a session, the state summary generated by the experiment controller also carries the current timestamp and remaining time in the session. Agents are expected to take the state summary as input and generate a response of actions, sometimes with rationales if they were asked to do so. The controller validates responses against the schemas and policies, returns an error message to the agent if an action was identified as invalid, and asks the agent to re-generate the response. Once the actions are validated, the controller commits these actions atomically, updates the environment state, and logs the whole interaction.
@@ 3.2.5. Experiment Controller
The experiment controller is the execution engine of the platform. During experiment setup, the controller creates a session instance, compiles experiment configurations, binds variables to the conditions specified by the researcher, and renders the user interfaces. During a live session, the controller manages the timer, validates actions and interactions for participants, updates private and shared state, and tracks all behavioral logs for post-hoc analysis. If an LLM agent is registered, the controller instantiates an agent context, as described above. At the same time, human participants can log in to the participant interface using their session ID and participant ID. The controller exhaustively tracks all the actions, including trades, messages, resource production and allocations, order submissions, as well as agent inputs and outputs.
3.3. Configurable Interaction Control for Collaboration
@@
This layer focuses on the type and visibility of information that is accessible to human participants and LLM agents. Prior research establishes that media characteristics shape mutual understanding (Clark and Brennan, 1991; Straus, 1997), awareness of others’ actions supports coordination (Gutwin and Greenberg, 2002; Dourish and Bellotti, 1992), and information control affects strategic behavior in distributed teams (Cramton, 2001; Olson and Olson, 2000). We implement two primary controls with respect to these conceptual findings. The Information Dashboard can be enabled or disabled to show real-time, shared information to all the participants, such as each other’s trades, recent orders, inventories, and order fulfillment status. Researchers can further customize information visibility based on participant groups (e.g., all, human-only, agent-only, etc.), change the information update frequency, and adjust the granularity of information (i.e., real values versus text summary). The Communication Transparency can be configured for strictly private (one-to-one), group-chat, or disabled entirely. This lets researchers examine how information sharing and message visibility influence negotiation dynamics and group behavior, and may further customize this control with additional turn-taking and message length limits.
@@
This layer defines the types of actions available, how they can be conducted, and the operational constraints that shape interdependence. Prior work links action coupling, protocols, and constraints to coordination cost, convergence, and coalition dynamics (Olson and Olson, 2000; Bos et al., 2004). The platform provides control over the Negotiation Protocol, which allows the trading interface to be configured for either strict accept-or-reject protocols or for open-ended negotiation with counter-offers. Researchers can also set constraints such as price limits or trade frequency. Researchers can also manage Concurrent Actions by either allowing participants to make multiple simultaneous trade offers, or participants must resolve one offer before making another
@@
This layer controls the timing and feedback mechanisms of LLM agents. Timing of responses can affect perceived competence and availability in remote collaboration (Olson and Olson, 2000), while adaptive communication and clear feedback help users recover from errors and calibrate reliance (Amershi et al., 2019). Key controls include Agent Response Latency, which can be set to be immediate or include a fixed or variable delay. Researchers can also customize a ”typing” or ”thinking” indicator. Agent Adaptive Feedback configures whether LLM agents use static response styles or dynamically adapt their strategy based on interaction history, specified through their persona profile. Finally, researchers have control over agents’ Explanations and Error Handling by deciding whether agents provide rationales proactively or on-demand. This control is extensible to support advanced agent reasoning architectures, such as the fast-and-slow thinking systems used in recent agent research (Lu et al., 2025).
@@ 3.4. Adaptivity Across Classic Experiments
The combination of the Experiment Configuration Language and the layered interaction controls allows the platform to be rapidly adapted for re-implementing classic experimental paradigms or creating novel ones. While our case studies in Section 4.1 provide an in-depth demonstration of a Shape Factory re-implementation (Bos et al., 2004), we describe here how the platform’s modularity supports several other paradigms from collaboration research. The participant interfaces correspond to each paradigm is shown in Fig 9.
One such paradigm is DayTrader (Bos et al., 2002), a social dilemma experiment that analyzes how communication media affect trust. In this task, participants make decisions about individual versus group investments over multiple rounds. To configure this experiment, the Trading page and My Task module are disabled, while the My Action module is adapted to support the investment decisions. The Information Dashboard is set to display group contributions and individual income, and the researcher enables the group chat in the Chat page to allow for discussion.
Another example is Essay Ranking (Zheng et al., 2023b), a group decision-making task originally conducted via a Wizard-of-Oz setup. Participants first review and rank essays independently, then discuss their rankings as a group before submitting a final independent ranking. Our platform supports this workflow by adapting the My Status module to display the essays and using the My Action module for the ranking inputs. The group discussion phase is enabled by allowing group chat in the Chat page.
Beyond these specific examples, the platform’s core components are sufficiently general to support a broad class of collaborative idea-generation and problem-solving tasks (Straus and McGrath, 1994). Such tasks, which range from brainstorming to solving logic problems, primarily depend on robust communication channels and the ability to present shared materials, both of which are native features of our platform. Similarly, two-player, turn-based paradigms like the ”Passcode” word-guessing game used to study mental models of AI (Gero et al., 2020) can be implemented by configuring the action structure for sequential, one-word message exchanges between two participants. This demonstrates the platform’s capacity to support diverse forms of collaborative interaction.
@@ 4. Evaluation of Our Platform
We conducted a two-part evaluation to assess two critical properties of our research platform. First, to measure control effectiveness, we evaluated the platform’s ability to support rigorous experiments that can produce significant and measurable findings. This is achieved through two controlled case studies using a re-implementation of the Shape Factory experiment (Bos et al., 2004) to test interaction controls. Second, to assess usability, we examined how easily target users could navigate the study preparation and analysis workflow. In particular, we conducted a participatory cognitive walkthrough with HCI researchers to evaluate the researcher interface and iterate the interface design.
@@ 4.1. Case Studies with Shape Factory Experiment
We selected the Shape Factory paradigm as the testbed for our case studies because it provides a versatile environment for examining complex collaborative dynamics. It involves open-ended negotiation, resource exchange, and multi-player communication under time pressure, and its core mechanics connect to numerous other classic CSCW experiments that isolate specific variables like trust (Bos et al., 2002) or role interdependence (McNeese et al., 2005; Schelble et al., 2022). This versatility makes Shape Factory an appropriate paradigm for validating whether our platform’s interaction controls can produce meaningful behavioral and perceptual differences. While our implementation adheres to the original experiment’s core rules, incentives, and resource scarcity design, our goal was not to replicate the original research topic and findings, but to validate our platform’s utility for human-agent collaboration research.

### v2

@@ 3.2.1. Participant Interface
The participant interface uses a three-column layout that can generalize across common experimental paradigms by separating visual modules for private state, social interaction, task operations, and shared awareness. An example participant interface for the Shape Factory experiment, which is used in the case study in Section 4.1, is shown in Figure 3.
The left column presents private state and individual actions. The My Status module shows individual resources and role-dependent capacities that are not necessarily visible to others. The My Action module provides affordances for private actions, such as producing a resource or making an independent decision. The My Task module displays individual or collaborative goals and fulfillment status. The right column is a fully configurable shared Information Dashboard. The dashboard makes workspace awareness (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002) a customizable visual module; researchers can select what information to disclose, to whom, and at what level of granularity.
The middle column contains the Social module for all participant interactions. This is supported by fruitful CSCW research that has underscored the importance of communication and negotiation affordances for studying trust, information grounding, and collaboration effectiveness (Straus, 1997; Bos et al., 2002; Hinds and Bailey, 2003). Here, a Trading page supports resource exchange and can be adaptive to broader collaborative decision-making activities. A separate Chat page supports text-based communication between participants. Researchers can configure trading with free-form or protocol-bounded offers to study negotiation dynamics and coalition formation (Poole et al., 1991). For the chat functionalities, we implemented three variations: private messages, group messages, or disabled them entirely.
@@ 3.2.2. Researcher Interface
The researcher interface provides direct controls for researchers to design, execute, monitor, and analyze experiments for human-agent collaboration research. The final design, as shown in Fig 4, was refined through a participatory cognitive walkthrough (Polson et al., 1992) with five HCI researchers by incorporating design and interaction feedback and aligning the workflow with their expectations (details in Section 4.3). The researcher interface organizes functionalities into three pages: experiment setup, real-time monitoring, and post-study analysis. For experiment setup, the interface guides researchers through four primary steps: experiment selection, parameter configuration, interaction control, and participant registration, as shown in Figure 4. In addition, researchers can customize participants’ information, including the persona profile for LLM agents, without editing the code.
The Experiment Selection module lets researchers load a saved session or create a new one. After a session is created, it presents a list of experimental paradigms with brief descriptions and simplified illustrations, allowing researchers to quickly understand each paradigm’s purpose even without prior familiarity. Researchers can also upload a custom paradigm defined in an ECL file.
Then, the Session Configuration module offers researchers comprehensive control over (1) experiment parameters (e.g., initial resource allocation and experiment constraints); (2) interaction controls, such as communication channels and information dashboard (details in Section 3.3); (3) agent design (e.g., response latency); (4) participant registration for both humans and agents.
Once the session is live, the Real-time Monitoring module allows researchers to observe participant interactions and progress. After the session is completed, the Result Analysis module provides summary statistics and visualizations of participant performance. Researchers can view aggregate or per-participant results and export raw event logs for further analysis.
Figure 4. The workflow of HCI researchers for study setup through the researcher interface, which includes five steps: experiment selection, parameter configuration, interaction control, participant registration, and participant customization.
ig. 4. Workflow for study setup in the researcher interface. The diagram outlines four steps for researchers: selecting an experiment, configuring parameters, controlling interactions, and registering participants.
@@
An overview of HCI researchers’ workflow is illustrated in Figure 4. In the study setup, researchers first decide whether to load an existing or a new session; if the latter, they must choose the experimental paradigms that best suit their research interests. The system also enables researchers to upload a customized experiment configuration, using ECL specified in Section 3.2.3, if the provided ones do not meet their needs. Then, researchers will proceed to the session configuration module to customize interaction controls and agent-specific designs. The horizontal scrolling function allows researchers to iteratively adjust parameters in accordance with experimental procedures. After completing all configurations, researchers conclude the setup by registering the human and agent participants.
@@ 3.2.3. Experiment Configuration Language (ECL)
The ECL is specified in a structured format by defining its components through a set of core primitives: objects, actions, policies, and views. The language is organized around these four fundamental concepts. Objects represent the primary entities within the experiment, such as resources (e.g., money, parts), roles (e.g., buyer, seller), or abstract concepts like team scores. Each object is defined by a set of attributes, such as initial_valueinitial\_value or visibilityvisibility. Actions define the operations available to participants, specifying the actor, associated costs, and their effects on the system’s object states. Policies are logical constraints that are specified by different experimental paradigms, ranging from preconditions for specific actions (e.g., requiring sufficient resources) to global rules like time limits or winning conditions. Finally, Views separates the system’s internal state from the presentation layer by defining which objects and attributes are rendered on the participant interface and how. This separation is critical, as it allows researchers to manipulate information visibility without altering the experiment’s core logic.
A concrete example of ECL for the Shape Factory is provided in Appendix B. In this example, $Money$ and $Shape$ are defined as objectsobjects. The $ProduceAction$ links the costs and effects for producing a shape, and is governed by a policy that validates the actor’s resources. The $Views$ declares what objects and states are visible in each visual module. During setup, the Experiment Controller parses an ECL file to dynamically instantiate the environment, enforce the specified policies, and render the correct views for participants. The experiment controller also validates the configurations and reports conflicts when a new configuration template is uploaded. This approach allows researchers to create, share, and modify experimental paradigm designs in a transparent and reproducible manner.
@@ 3.2.4. Agent Context Protocol
Fig. 5. Agent Context Protocol (ACP). This schematic depicts ACP as a unified communication channel that connects the experiment controller with large language model agents, ensuring standardized integration regardless of the agent’s internal design.
Inspired by the emerging standard of model context protocol [^3] https://modelcontextprotocol.io/ that connects LLMs with external tools and resources, we designed the ACP to standardize communication between LLM agents and the experiment environment (Figure 5). ACP aims to ensure agents’ behaviors align with the experimental rules and valid action spaces, regardless of agent architecture. More importantly, agents perceive and act under the same constraints and through the same structured channels as remote human participants, ensuring parity for controlled, comparative studies.
At initialization, the experiment controller (Section 3.2.5) constructs an experiment-specific context for agents from the experiment configurations and the interaction conditions specified by researchers. The context includes participant IDs, groups, private and public states, permitted actions with schemas, communication channels, perception interval, and experiment rules. The perception interval determines how often the controller sends state summaries to agents. The controller also validates that each summary is consistent with the pre-defined private and public states for the agent.
During a session, each state summary from the controller includes the current timestamp and remaining session time. Agents are expected to take the state summary as input and generate a response of actions, sometimes with rationales if required. The controller validates responses against the schemas and policies, returns an error message to the agent if an action was identified as invalid, and asks the agent to re-generate the response. Once the actions are validated, the controller commits these actions atomically, updates the environment state, and logs the whole interaction.
@@ 3.2.5. Experiment Controller
The experiment controller is the execution engine of the platform. During experiment setup, the controller creates a session instance, compiles experiment configurations, binds variables to researcher-specified conditions, and renders the user interfaces. During a live session, the controller manages timing, validates participant actions and interactions, updates private and shared states, and records behavioral logs for post-hoc analysis. If an LLM agent is registered, the controller instantiates an agent context, as described above. Meanwhile, human participants can log in to the participant interface using their session ID and participant ID. The controller exhaustively tracks all actions, including trades, messages, resource production and allocation, order submissions, and agent inputs and outputs.
To safeguard participant data, the controller stores logs locally by default, avoiding third-party cloud retention. To prevent model ’jailbreaks’ or toxic output, the ACP enforces strict schema validation on all agent outputs before rendering them to humans, which can effectively sandbox the agent’s communicative capabilities.
3.3. Configurable Interaction Control for Human-Agent Collaboration
@@
This layer focuses on the type and visibility of information that is accessible to human participants and LLM agents. Prior research establishes that media characteristics shape mutual understanding (Clark and Brennan, 1991; Straus, 1997), awareness of others’ actions supports coordination (Gutwin and Greenberg, 2002; Dourish and Bellotti, 1992), and information control affects strategic behavior in distributed teams (Cramton, 2001; Olson and Olson, 2000). We implement two primary controls with respect to these conceptual findings. The Information Dashboard can be enabled or disabled to show real-time, shared information to all the participants, such as each other’s trades, recent orders, inventories, and order fulfillment status. Researchers can further customize information visibility based on participant groups (e.g., all, human-only, agent-only, etc.), change the information update frequency, and adjust the granularity of information (i.e., real values versus text summary). The Communication Transparency can be configured for strictly private (one-to-one), group chat, or disabled entirely. This lets researchers examine how information sharing and message visibility influence negotiation dynamics and group behavior, and may further customize this control with additional turn-taking and message length.
@@
This layer defines the types of actions available, how they can be conducted, and the operational constraints that shape interdependence. Prior work links action coupling, protocols, and constraints to coordination cost, convergence, and coalition dynamics (Olson and Olson, 2000; Bos et al., 2004). The platform provides control over the Negotiation Protocol, which allows either strict accept-or-reject protocols or open-ended negotiation with counter-offers. Researchers can also set constraints such as price limits or trade frequency. Researchers can also manage Concurrent Actions by either allowing participants to make multiple offers simultaneously, or participants must resolve one before making another.
@@
This layer controls the timing and feedback mechanisms of LLM agents. Timing of responses can affect perceived competence and availability in remote collaboration (Olson and Olson, 2000), while adaptive communication and clear feedback help users recover from errors and calibrate reliance (Amershi et al., 2019). Key controls include Agent Response Latency, which can be set to be immediate or include a fixed or variable delay. Researchers can also customize a ”typing” or ”thinking” indicator. Agent Adaptive Feedback configures whether LLM agents use static response styles or dynamically adapt their strategy based on interaction history, specified through their persona profile. Finally, researchers have control over agents’ Explanations and Error Handling by deciding whether agents provide rationales proactively or on demand. This control is extensible to support advanced agent reasoning architectures, such as the fast-and-slow thinking systems used in recent agent research (Lu et al., 2025).
@@ 3.4. Adaptivity Across Classic Experiments
The combination of the Experiment Configuration Language and the layered interaction controls allows the platform to be rapidly adapted for re-implementing classic experimental paradigms or creating novel ones. While our case studies in Section 4.1 provide an in-depth demonstration of a Shape Factory re-implementation (Bos et al., 2004), we describe here how the platform’s modularity supports several other paradigms from collaboration research. The participant interfaces corresponding to each paradigm are shown in Fig 9.
One such paradigm is DayTrader (Bos et al., 2002), a social dilemma experiment that analyzes how communication media affect trust. In this task, participants make decisions about individual versus group investments across multiple rounds. To configure this experiment, the Trading page and My Task module are disabled, while the My Action module is adapted to support investment decisions. The Information Dashboard displays group contributions and individual income, and group chat in the Chat page is enabled for discussion.
Another example is Essay Ranking (Zheng et al., 2023a), a collaborative decision-making task originally conducted via a Wizard-of-Oz setup. Participants first review and rank essays independently, then discuss their rankings as a group before submitting a final independent ranking. Our platform supports this workflow by adapting the My Status module to display the essays and using the My Action module for the ranking inputs. The group discussion phase is enabled by allowing group chat in the Chat page.
Beyond these specific examples, the platform’s core components are sufficiently generalizable to support a broad class of collaborative idea-generation and problem-solving tasks (Straus and McGrath, 1994). Such tasks, which range from brainstorming to solving logic problems, primarily depend on robust communication channels and the ability to present shared materials, both of which are native features of our platform. Similarly, two-player, turn-based paradigms like the ”Passcode” word-guessing game for studying humans’ mental models of AI (Gero et al., 2020) can be implemented by configuring the action structure for sequential, one-word message exchanges between two participants.
@@ 4. Evaluation of Our Platform
We conducted a multi-phase evaluation to assess the platform’s research efficacy as an instrument for controlled studies, its extensibility across collaboration paradigms, and its usability for researchers. Specifically, we conducted two controlled experiments with human participants to examine the platform’s ability to effectively manipulate theory-driven interaction controls and test whether the manipulations yield measurable changes in outcomes and perceptions. The first study utilized the Shape Factory (Bos et al., 2004) paradigm, a resource-negotiation task, with the manipulation of the Information Flow controls (Section 4.1). The second study utilized the Hidden Profile (Stasser and Titus, 1985, 1987) paradigm, a collaborative problem-solving task, by manipulating the Agent Initiative controls (Section 4.2). Finally, we conducted a participatory cognitive walkthrough with HCI researchers of diverse experience to examine the researcher workflow and interface design (Section 4.3).
@@ 4.1. Case Studies with Shape Factory Experiment
We selected the Shape Factory paradigm as the first testbed for our case studies because it provides a versatile environment for examining complex collaborative dynamics. It involves open-ended negotiation, resource exchange, and multi-player communication under time pressure, and its core mechanics connect to numerous other classic CSCW experiments that isolate specific variables like trust (Bos et al., 2002) or role interdependence (McNeese et al., 2005; Schelble et al., 2022). While our implementation adheres to the original experiment’s core rules, incentives, and resource scarcity design, our goal was not to replicate the original research topic and findings, but to validate our platform’s utility for human-agent collaboration research.